OpenStudy (owen_turner):

What is the point-slope form of the equation of a line that passes through the point (1, 4) and has a slope of −3? A. y-1=-3(x-4) B. y+4=3(x+1) C. y+4=-3(x+1) D. y-4=-3(x-1)

OpenStudy (owen_turner):

Help? Anyone?

OpenStudy (sooobored):

point-slope form \[y-y_1=m(x-x_1)\] m is generally considered the slope so what's left is determine which is x1 and which is y1 from (1,4)

OpenStudy (owen_turner):

So it would be D.?

Directrix (directrix):

D. y-4=-3(x-1) That is correct. @Owen_Turner